Two equal positive point charges q1=q2=2µC interact with a third point charge q3=4µC. Find

the magnitude and direction of the total (net) force on q3. (q1 is 0.3m above the origin; q2 is

0.3m below the origin; q3 is 0.4m right from the origin).


1
Expert's answer
2021-01-30T16:47:59-0500

Let's write the forces that act on charge "q_3" due to charges "q_1"and "q_2":


"F_1=\dfrac{kq_1q_3}{r^2},""F_2=\dfrac{kq_2q_3}{r^2}."


We can find the distance from the charges "q_1"and "q_3" as well as "q_2" and "q_3" from the Pythagorean theorem:


"r=\sqrt{(0.3\ m)^2+(0.4\ m)^2}=0.5\ m."

Since, "q_1=q_2" and "r_1=r_2=r" the magnitudes of two forces are equal. Let's write the components of forces "F_1" and "F_2" in projections on "x" and "y" axis:


"F_1cos\theta+F_2cos\theta=2F_1cos\theta,""-F_1sin\theta+F_2sin\theta=0."

As we can see, "y"-components cancel out, so we can write the net force on "q_3":


"F_{net}=2\dfrac{kq_1q_3}{r^2}cos\theta."

We can find angle "\theta" from the geometry:


"\theta=cos^{-1}(\dfrac{0.4\ m}{0.5\ m})=36.86^{\circ}."

Finally, we can find the net force on "q_3":


"F_{net}=2\cdot\dfrac{9\cdot10^9\ \dfrac{Nm^2}{C^2}\cdot(2\cdot10^{-6}\ C)^2}{(0.5\ m)^2}cos36.86^{\circ}=0.23\ N."

Answer:

"F_{net}=0.23\ N," in "+x"-direction.